19 May 2017
Friday 19th of May 2017

Iben is loading ice core troughs into the core storage.

The Twin Otter made it to Summit today.

This morning weather improved, and the Twin Otter was able to fly to Summit, and that reduced camp population from 35 to 29. All the while work continued on many fronts.

In the drill trench, the drillers are mounting the last of the infrastructure needed to begin drilling with drilling fluid. Jan from company SVEDAN is mounting the freezer system to keep the ice cores cold while extracting them from the drill.

Elizabeth and Nanna are working on a system on probes that allow for wireless transfer of data from inside the ice sheet. Harald is setting up a system to monitor the atmospheric boundary layer.

Although it is considerably warmer than last week, wind is still an issue and this kept us from completing WP 2. It will be finished tomorrow. Inside the trenches there is activity in both science trench and drill trench. While the drill trench is almost complete, the science trench is slowly being outfitted with table saws and equipment.

What we did today:

  1. GL 890 Twin Otter managed to fly to Summit.
  2. Removing snow around carpenters garage.
  3. Removing snow drifts from entrance to trenches.
  4. Installing slush pans, railing and drill fluid gutter at drill site.
  5. Construction of WP 2 almost completed.
  6. Work on installing the active cooling system for the ice core extraction freezer.
  7. German snow sampling.
  8. Japanese aerosol sampling and snow experiments.
  9. Water vapour sampling, methane sampling, boundary layer and meteorology station working.
  10. Work on the “egg” project.
  11. Work on setup of system to measure atmospheric boundary layer dynamics.
  12. 2.5 m high resolution snow pit sampling completed.

Weather today: Sunshine with thin high clouds. Temp. - 6°C to - 20°C. Wind: 8 - 23 kt from SW. Visibility: To horizon.

← Previous entry   Next entry →